# GENPACK: KPI-Guided Multi-Objective Genetic Algorithm for Industrial 3D Bin Packing

### v2

@@ Abstract.
The three-dimensional bin packing problem (3D-BPP) is a longstanding challenge in operations research and logistics. Classical heuristics and constructive methods can generate packings quickly, but often fail to address industrial constraints such as stability, balance, and handling feasibility. Metaheuristics such as genetic algorithms (GAs) provide flexibility and the ability to optimize across multiple objectives; however, pure GA approaches frequently struggle with efficiency, parameter sensitivity, and scalability to industrial order sizes. This gap is especially evident when scaling to real-world pallet dimensions, where even state-of-the-art algorithms often fail to achieve robust, deployable solutions. We propose a KPI-driven GA-based pipeline for industrial 3D-BPP that integrates key performance indicators (KPIs) directly into a multi-objective fitness function. The methodology combines a layer-based chromosome representation with domain-specific operators and constructive heuristics to balance efficiency and feasibility. On the BED-BPP benchmark of 1,500 real-world orders, our Hybrid-GA pipeline consistently outperforms heuristic- and learning-based state-of-the-art methods, achieving up to 35% higher space utilization and 15–20% stronger surface support, with lower variance across orders. These improvements come at modest runtime cost but remain feasible for batch-scale deployment, yielding stable, balanced, and space-efficient packings.
genetic algorithms, 3D bin packing, multi-objective optimization, industrial logistics, metaheuristics, chromosome encoding
@@ 1. Introduction
The three-dimensional bin packing problem (3D-BPP) is a fundamental NP-hard optimization challenge with significant implications for logistics, warehouse management, and supply chain operations. The objective is to arrange a set of cuboidal items into a limited number of bins or pallets while minimizing void volume and satisfying geometric feasibility (martello2000three, ). Classical formulations and early heuristics primarily optimize for volume utilization or bin count, but such objectives alone are insufficient for industrial palletizing. In practice, additional requirements such as stability, support, handling feasibility, and weight distribution must be satisfied for solutions to be deployable (bortfeldt2013constraints, ).
@@
Constructive heuristics such as Bottom-Left (chazelle1983bottomleft, ), Extreme Point (crainic2008extreme, ), or MaxRects (jylanki2010maxrects, ) remain widely used for their efficiency, particularly on large instances (bischoff1990three, ). However, they rely heavily on simplified placement rules that often yield arrangements failing to satisfy industrial safety and stability constraints. Metaheuristics, and in particular genetic algorithms (GAs), provide greater flexibility by exploring large solution spaces and accommodating multiple objectives (jakobs1996genetic, ; wu2010three, ). However, pure GA approaches also face limitations: common encodings are not well suited to the packing structure, genetic operators often fail to preserve feasibility, and fitness functions often fail to incorporate practical indicators of industrial quality (bortfeldt2001hybrid, ; egeblad2009heuristic, ; bortfeldt2013constraints, ). Moreover, scalability to real-world order sizes at industrial pallet dimensions, such as Euro-pallets and roll containers, remains a critical challenge where many state-of-the-art methods still struggle (wu2010three, ; egeblad2009heuristic, ).
@@
Recent work has introduced hybrid methods that combine heuristics with metaheuristic refinement. For example, Ananno and Ribeiro (ananno2024multiheuristic, ) proposed a two-stage framework in which constructive heuristics generate layers and blocks, and a GA arranges the remaining residual items under stability and support constraints. While such approaches demonstrate feasibility, they primarily enforce industrial requirements as hard constraints and focus evaluation on compactness and utilization (bortfeldt2001hybrid, ; kang2012hybrid, ; egeblad2009heuristic, ). They do not systematically optimize across multiple industrial key performance indicators (KPIs), and runtime scalability remains problematic for highly heterogeneous orders (egeblad2009heuristic, ; wu2010three, ).
@@
This paper addresses the above limitations by presenting a KPI-driven GA specifically designed for industrial 3D-BPP. The overall three-phase pipeline of the proposed approach is shown in Figure 1. In the first phase, constructive heuristics such as MaxRects (jylanki2010maxrects, ) rapidly generate initial placements that only cover a subset of all the items to be packed. The second phase applies a genetic algorithm, where selection, crossover, mutation, and KPI-based fitness evaluation are used to refine the packing and place residual items that the heuristics failed to place. Finally, a post-processing step compacts and adjusts the arrangement to maximize density and ensure feasibility with respect to stability and industrial constraints.
@@
Figure 1. Three-phase pipeline architecture: heuristics generate initial placements, a genetic algorithm refines residuals, and post-processing optimizes the final configuration.
@@
Diagram showing a three-phase pipeline: heuristics for initial placements, genetic algorithm for refinement, and post-processing for optimization.
@@
  • A multi-objective fitness formulation integrating industrially relevant KPIs such as density, stability, surface support, balance, and structural integrity.
@@
The remainder of the paper is structured as follows: Section 2 reviews related work, Section 3 defines the problem formulation, Section 4 presents the proposed GA methodology, Section 5 describes the hybrid pipeline, Section 6 discusses implementation and dataset characteristics, Section 7 reports experimental results, Section 8 provides a broader discussion, and Section 9 concludes with future directions.
@@ 2. Related Work
Research on the 3D-BPP spans heuristics, metaheuristics, and learning methods, alongside more recent, industrially motivated formulations.
@@ 2.1. Heuristics
Constructive heuristics such as Bottom-Left (chazelle1983bottomleft, ), Extreme Point (crainic2008extreme, ), and MaxRects (jylanki2010maxrects, ) generate feasible layouts quickly and scale to large instances (bischoff1990three, ). Their efficiency makes them useful in practice, but they usually rely on simple placement rules that often result in unstable or unbalanced packings, which usually cannot be compensated by post-processing to meet industrial constraints.
@@ 2.3. Learning-based Approaches
Learning-based methods replace explicit search with learned inference and deliver fast test-time decisions, but they often require large training corpora and careful reward and constraint design, and they can be sensitive to domain shift between synthetic training distributions and industrial orders. Adapting to specific operational rules or KPI trade-offs often necessitates retraining, and additional post-processing may still be required for feasibility on out-of-distribution cases (zhao2021constrained, ; zhao2022pct, ; que2023transformer, ; xiong2024gopt, ).
@@ 2.5. Benchmarking
Traditional benchmarking has largely relied on synthetic datasets such as the Bischoff and Marriott instances (bischoff1990three, ) and, more recently, the software toolbox for dataset generation (ribeiro2023toolbox, ). While useful for reproducibility, these datasets are either fully synthetic or designed primarily for multi-container problems, and thus fail to capture the heterogeneity and operational constraints of single-container, real-world palletization scenarios. In contrast, the BED-BPP dataset (kagerer2023bedbpp, ), with more than 10,000 grocery orders including precise item dimensions and corresponding weights, enables realistic, industry-grounded evaluation and has therefore been adopted for empirical analysis in our work.
The above studies demonstrate the efficiency of heuristics in generating fast and interpretable packing solutions, the flexibility of metaheuristics in exploring diverse search spaces and refining constructive outputs, and the promise of learning-based policies in capturing structural patterns. At the same time, they also reveal limitations, such as simplified objective formulations that overlook practical KPIs like stability, load balance, or number of pallets, as well as runtime bottlenecks when scaling to real-world pallet dimensions. These gaps collectively motivate the design of a KPI-driven hybrid genetic algorithm in this work, aiming to maximize computational efficiency with practical relevance at industrial scale.
@@ 3.1. Setting and Variables
Given is a set of items $\mathcal{I}=\{1,\dots,n\}$, where item $i$ has dimensions $(l_{i},w_{i},h_{i})$ and mass $m_{i}$. Bins (pallets) share base $(L,W)$, height $H$, and payload capacity $M_{max}$. Items are assumed to be axis-aligned in a right-handed coordinate system where $z$ is normal to the pallet base. Rotation is restricted to $90^{\circ}$ multiples around the $z$-axis, which allows interchanging the $x$- and $y$-dimensions of an item while keeping its height upright. This reflects common industrial requirements, where most packaged goods can be safely turned sideways on the pallet but should not be tilted onto their ends, as such orientations risk compromising structural integrity or violating handling conventions (bortfeldt2013constraints, ). A packing solution assigns each item to one bin with a placement $(x_{i},y_{i},z_{i})$, denoting the coordinates of its lower–left–front corner, i.e., the corner at minimal $(x,y,z)$ in the global reference frame, together with a rotation $\theta_{i}$.
@@ 3.2. Feasibility
These constraints are enforced during both construction and refinement, for example, by axis-disjunctive checks to prevent overlaps and per-bin mass accumulation to enforce payload limits (martello2000three, ). Additional rules, such as minimal support or restricted orientations, can be activated depending on the specific problem instance (bortfeldt2013constraints, ).
@@ 4.1. Chromosome Representation
where each layer $\mathcal{L}_{j}$ stores (i) its assigned items, (ii) their 2D coordinates $\{(x_{i},y_{i})\}$, (iii) its height $h(\mathcal{L}_{j})$, the maximum height among contained items, and (iv) the layer’s base $z$-level. Layers are stacked sequentially, and the cumulative height defines overall packing height. Within a layer, items are axis-aligned and placed on a coarse grid that avoids 2D overlaps and vertical overlaps are prevented by stacking order.
@@ 4.2. Multi-KPI Fitness Function
A single scalar fitness aggregates multiple KPIs (7 within this work) and an overlap penalty. For chromosome $C$:
@@
The KPIs in Eq. (3) combine standard measures of volume utilization, support, and center-of-gravity balance, which are widely used in container loading and 3D-BPP studies (martello2000three, ; bortfeldt2013constraints, ; crainic2008extreme, ; ramos2018new, ; elhedhli2019three, ), with problem-specific extensions (e.g., RelDen, TallItem) that we introduce to capture compactness and stability in pallet loading more effectively. They are defined as:
@@
  • RelDen: relative density below the upper envelope, capturing void reduction and local compactness. While compaction is a common heuristic principle in 3D-BPP algorithms (crainic2008extreme, ; egeblad2009heuristic, ), this envelope-based voxel KPI is specific to this work. Using a voxel grid of spatial resolution $r$ discretizing the pallet base $(L,W)$ and height $H^{\ast}$, let $\mathcal{G}[x,y,z]\in\{0,1\}$ denote occupancy (1 if voxel is filled). Define the upper envelope
@@
  • SideSup: lateral support. Stability requirements that restrict unsupported side faces are common in container and pallet loading (bortfeldt2013constraints, ). For each item, compute the area fraction of each vertical face that is in contact with neighboring items or pallet walls; a face is considered supported if at least $20\%$ of its area is covered. SideSup is the average supported-face rate over all items.
@@
    where $\delta$ denotes the normalized radial distance (Euclidean offset) of the total 2D center of gravity from the pallet center and
@@
All KPI thresholds (e.g., the $20\%$ side-contact in SideSup and $75\%$ base-overlap in SurfSup) were empirically determined through ablation studies and can be adjusted for different problem settings. They are consistent with established practices in container and pallet loading (bortfeldt2013constraints, ).
@@ 4.3. Population Initialization
The initial population $\mathbf{P}_{0}$ is built from residual superitems $R$ that were not included in the constructive phase solution $S_{0}$, i.e., items left unpacked during the heuristic phase. It follows the procedure in Algorithm  1 combining multiple low-cost heuristics to ensure diversity. Specifically, items are sorted by volume (VolSort), by base area (AreaSort), or left in random order (Random); spatial placement is then performed layer-wise on a coarse grid using the BottomLeft rule (chazelle1983bottomleft, ; crainic2008extreme, ) or an ExtremePoint (EP) strategy (crainic2008extreme, ).
@@ 4.4. Evolutionary Operators
4.4.1. Selection and Replacement:
We use a tournament selection with adaptive pressure defined as
@@ 4.4.1. Selection and Replacement:
where POP denotes the population size, $g$ the current generation, and $G$ the total number of generations. This formulation gradually increases the tournament size, from roughly 10% to 40% of the population. The best 10% of individuals (elitism) are preserved in each generation, and the remaining offspring are produced from the selected parents.
4.4.2. Crossover and Mutation:
A one-point crossover with layer preservation is applied with probability:
@@ 4.4.2. Crossover and Mutation:
exchanging layer suffixes between parents and recomputing $z$-levels and heights. If feasibility cannot be maintained, parents are copied unchanged.
Mutation occurs with probability:
@@
favoring exploration early in evolution. Each mutation operator $\text{MUT}_{k}\in\{\text{MUT}_{k}\}_{k}$ corresponds to a KPI-aligned adjustment. Specifically, mutations include shifting tall items toward the center or lower layers to improve TallItem and CoG2D, filling small 2D voids below items to improve RelDen, compacting arrangements toward the origin to enhance AbsDen and RelDen, optimizing lateral contact to improve SideSup, increasing bottom overlap to strengthen SurfSup. Additional diversification operations such as item, layer, or layer-addition swaps are also applied.

### v3

# GENPACK: KPI-Guided Multi-Criteria Genetic Algorithm for Industrial 3D Bin Packing

@@ Abstract.
The three-dimensional bin packing problem (3D-BPP) is a longstanding challenge in operations research and logistics. While classical heuristics and constructive methods can generate packings efficiently, they often fail to satisfy industrial requirements such as stability, balance, and handling feasibility. Metaheuristics such as genetic algorithms (GAs) offer greater flexibility, but pure GA approaches frequently struggle with efficiency, parameter sensitivity, and scalability to industrial order sizes. These limitations are particularly evident at real-world pallet dimensions, where even state-of-the-art methods often fail to produce robust, deployable solutions. We propose a KPI-guided GA-based pipeline for industrial 3D-BPP that integrates key performance indicators (KPIs) directly into a scalarized fitness function. The method combines a layer-based chromosome representation, domain-specific operators, and constructive heuristics to balance efficiency and feasibility. On the BED-BPP benchmark of 1,500 real-world orders, our GENPACK pipeline consistently outperforms heuristic and learning-based baselines, achieving up to 35% higher space utilization and 15–20% stronger surface support, while exhibiting lower variance across orders. These gains come at a modest runtime cost but remain practical for batch-scale deployment, yielding stable, balanced, and space-efficient packings.
genetic algorithms, 3D bin packing, multi-criteria optimization, industrial logistics, metaheuristics, chromosome encoding
@@ 1. Introduction
The three-dimensional bin packing problem (3D-BPP) is a fundamental NP-hard optimization challenge with direct relevance to logistics, warehouse management, and supply chain operations. The objective is to arrange a set of cuboidal items into a limited number of bins or pallets while minimizing void volume and satisfying geometric feasibility (martello2000three, ). Classical formulations and early heuristics primarily optimize for volume utilization or bin count, but such objectives alone are insufficient for industrial palletizing. In practice, additional requirements such as stability, support, handling feasibility, and weight distribution must also be satisfied for solutions to be deployable (bortfeldt2013constraints, ).
@@
Constructive heuristics such as Bottom-Left (chazelle1983bottomleft, ), Extreme Point (crainic2008extreme, ), and MaxRects (jylanki2010maxrects, ) remain widely used because of their efficiency, particularly on large instances (bischoff1990three, ). However, they rely on simplified placement rules that often yield arrangements that fail to satisfy industrial safety and stability constraints. Metaheuristics, and genetic algorithms (GAs) in particular, provide greater flexibility by exploring larger solution spaces and accommodating multiple objectives (jakobs1996genetic, ; wu2010three, ). However, pure GA approaches also face important limitations: common encodings are not well suited to packing structure, genetic operators often fail to preserve feasibility, and fitness functions frequently do not capture practical indicators of industrial quality (bortfeldt2001hybrid, ; egeblad2009heuristic, ; bortfeldt2013constraints, ). Moreover, scalability to real-world order sizes at industrial pallet dimensions, such as Euro-pallets and roll containers, remains a critical challenge where many state-of-the-art methods still struggle (wu2010three, ; egeblad2009heuristic, ).
@@
Recent work has introduced hybrid methods that combine constructive heuristics with metaheuristic refinement. For example, Ananno and Ribeiro (ananno2024multiheuristic, ) proposed a two-stage framework in which constructive heuristics generate layers and blocks, and a GA arranges the remaining residual items under stability and support constraints. While such approaches demonstrate feasibility, they primarily enforce industrial requirements as hard constraints and focus evaluation on compactness and utilization (bortfeldt2001hybrid, ; kang2012hybrid, ; egeblad2009heuristic, ). They do not systematically optimize across multiple industrial key performance indicators (KPIs), and runtime scalability remains problematic for highly heterogeneous orders (egeblad2009heuristic, ; wu2010three, ).
@@
This paper addresses these limitations by presenting a KPI-guided GA specifically designed for industrial 3D-BPP. The overall three-stage pipeline of the proposed approach is shown in Figure 1. In the first stage, constructive heuristics such as MaxRects (jylanki2010maxrects, ) rapidly generate initial placements that cover only a subset of the items to be packed. In the second stage, a genetic algorithm uses selection, crossover, mutation, and KPI-based fitness evaluation to refine the packing and place residual items that the heuristics failed to place. In the final stage, post-processing compacts and adjusts the arrangement to improve density and ensure feasibility with respect to stability and industrial constraints.
@@
Figure 1. Three-stage pipeline architecture: constructive heuristics generate initial placements, a genetic algorithm refines residuals, and post-processing improves the final configuration.
@@
Diagram showing a three-stage pipeline: constructive heuristics for initial placements, a genetic algorithm for refinement, and post-processing for optimization.
@@
  • A scalarized KPI-guided fitness formulation integrating industrially relevant KPIs such as density, stability, surface support, balance, and structural integrity.
@@
The remainder of the paper is structured as follows. Section 2 reviews related work, Section 3 defines the problem formulation, Section 4 presents the proposed GA methodology, Section 5 describes the hybrid pipeline, Section 6 discusses implementation and dataset characteristics, Section 7 reports experimental results, Section 8 provides a broader discussion, and Section 9 concludes with future directions.
@@ 2. Related Work
Research on the 3D-BPP spans heuristics, metaheuristics, and learning-based methods, alongside more recent industrially motivated formulations.
@@ 2.1. Heuristics
Constructive heuristics such as Bottom-Left (chazelle1983bottomleft, ), Extreme Point (crainic2008extreme, ), and MaxRects (jylanki2010maxrects, ) generate feasible layouts quickly and scale to large instances (bischoff1990three, ). Their efficiency makes them useful in practice, but they usually rely on simple placement rules that often result in unstable or unbalanced packings, which typically cannot be sufficiently corrected by post-processing to meet industrial constraints.
@@ 2.3. Learning-based Approaches
Learning-based methods replace explicit search with learned inference and deliver fast test-time decisions, but they often require large training corpora and careful reward and constraint design, and they can be sensitive to domain shift between synthetic training distributions and industrial orders. Adapting them to specific operational rules or KPI trade-offs often necessitates retraining, and additional post-processing may still be required for feasibility on out-of-distribution cases (zhao2021constrained, ; zhao2022pct, ; que2023transformer, ; xiong2024gopt, ).
@@ 2.5. Benchmarking
Traditional benchmarking has largely relied on synthetic datasets such as the Bischoff and Marriott instances (bischoff1990three, ) and, more recently, software toolboxes for dataset generation (ribeiro2023toolbox, ). While useful for reproducibility, these datasets are either fully synthetic or designed primarily for multi-container problems, and thus fail to capture the heterogeneity and operational constraints of single-container, real-world palletization scenarios. In contrast, the BED-BPP dataset (kagerer2023bedbpp, ), with more than 10,000 grocery orders including precise item dimensions and corresponding weights, enables realistic, industry-grounded evaluation and has therefore been adopted for empirical analysis in our work.
The above studies demonstrate the efficiency of heuristics in generating fast and interpretable packing solutions, the flexibility of metaheuristics in exploring diverse search spaces and refining constructive outputs, and the promise of learning-based policies in capturing structural patterns. At the same time, they also reveal limitations, such as simplified objective formulations that overlook practical KPIs like stability, load balance, or number of pallets, as well as runtime bottlenecks when scaling to real-world pallet dimensions. These gaps collectively motivate the design of a KPI-guided hybrid genetic algorithm in this work, aiming to maximize computational efficiency while maintaining practical relevance at industrial scale.
@@ 3.1. Setting and Variables
Given is a set of items $\mathcal{I}=\{1,\dots,n\}$, where item $i$ has dimensions $(l_{i},w_{i},h_{i})$ and mass $m_{i}$. Bins (pallets) share base $(L,W)$, height $H$, and payload capacity $M_{max}$. Items are assumed to be axis-aligned in a right-handed coordinate system where $z$ is normal to the pallet base. Rotation is restricted to $90^{\circ}$ multiples around the $z$-axis, which allows interchanging the $x$- and $y$-dimensions of an item while keeping its height upright. This reflects common industrial requirements, where most packaged goods can be safely turned sideways on the pallet but should not be tilted onto their ends, as such orientations risk compromising structural integrity or violating handling conventions (bortfeldt2013constraints, ). A packing solution assigns each item to one bin with a placement $(x_{i},y_{i},z_{i})$, denoting the coordinates of its lower-left-front corner, i.e., the corner at minimal $(x,y,z)$ in the global reference frame, together with a rotation $\theta_{i}$.
@@ 3.2. Feasibility
These constraints are enforced during both construction and refinement, for example by axis-disjunctive checks to prevent overlaps and per-bin mass accumulation to enforce payload limits (martello2000three, ). Additional rules, such as minimum support or restricted orientations, can be activated depending on the specific problem instance (bortfeldt2013constraints, ).
@@ 4.1. Chromosome Representation
where each layer $\mathcal{L}_{j}$ stores (i) its assigned items, (ii) their 2D coordinates $\{(x_{i},y_{i})\}$, (iii) its height $h(\mathcal{L}_{j})$, i.e., the maximum height among contained items, and (iv) the layer’s base $z$-level. Layers are stacked sequentially, and the cumulative height defines the overall packing height. Within a layer, items are axis-aligned and placed on a coarse grid that avoids 2D overlaps, while vertical overlaps are prevented by stacking order.
@@ 4.2. Multi-KPI Fitness Function
A single scalar fitness aggregates multiple KPIs (seven in this work) and an overlap penalty. For chromosome $C$:
@@
The KPIs in Eq. (3) combine standard measures of volume utilization, support, and center-of-gravity balance, which are widely used in container loading and 3D-BPP studies (martello2000three, ; bortfeldt2013constraints, ; crainic2008extreme, ; ramos2018new, ; elhedhli2019three, ), with problem-specific extensions (e.g., RelDen, TallItem) that we introduce to capture compactness and stability in pallet loading more effectively. They are defined as follows:
@@
  • RelDen: relative density below the upper envelope, capturing void reduction and local compactness. While compaction is a common heuristic principle in 3D-BPP algorithms (crainic2008extreme, ; egeblad2009heuristic, ), this envelope-based voxel KPI is specific to this work. Using a voxel grid of spatial resolution $r$ discretizing the pallet base $(L,W)$ and height $H^{\ast}$, let $\mathcal{G}[x,y,z]\in\{0,1\}$ denote occupancy (1 if a voxel is filled). Define the upper envelope
@@
  • SideSup: lateral support. Stability requirements that restrict unsupported side faces are common in container and pallet loading (bortfeldt2013constraints, ). For each item, we compute the area fraction of each vertical face that is in contact with neighboring items or pallet walls; a face is considered supported if at least $20\%$ of its area is covered. SideSup is the average supported-face rate over all items.
@@
    where $\delta$ denotes the normalized radial distance (Euclidean offset) of the total 2D center of gravity from the pallet center, and
@@
All KPI thresholds (e.g., the $20\%$ side-contact threshold in SideSup and the $75\%$ base-overlap threshold in SurfSup) were empirically determined through ablation studies and can be adjusted for different problem settings. They are consistent with established practices in container and pallet loading (bortfeldt2013constraints, ).
@@ 4.3. Population Initialization
The initial population $\mathbf{P}_{0}$ is built from residual superitems $R$ that were not included in the constructive phase solution $S_{0}$, i.e., items left unpacked during the heuristic phase. It follows the procedure in Algorithm 1, combining multiple low-cost heuristics to ensure diversity. Specifically, items are sorted by volume (VolSort), by base area (AreaSort), or left in random order (Random); spatial placement is then performed layer-wise on a coarse grid using the BottomLeft rule (chazelle1983bottomleft, ; crainic2008extreme, ) or an ExtremePoint (EP) strategy (crainic2008extreme, ).
@@ 4.4. Evolutionary Operators
4.4.1. Selection and Replacement
We use tournament selection with adaptive pressure defined as
@@ 4.4.1. Selection and Replacement
where POP denotes the population size, $g$ the current generation, and $G$ the total number of generations. This formulation gradually increases the tournament size from roughly 10% to 40% of the population. The best 10% of individuals (elitism) are preserved in each generation, and the remaining offspring are produced from the selected parents.
4.4.2. Crossover and Mutation
A one-point crossover with layer preservation is applied with probability
@@ 4.4.2. Crossover and Mutation
exchanging layer suffixes between parents and recomputing $z$-levels and heights. If feasibility cannot be maintained, the parents are copied unchanged.
Mutation occurs with probability
@@
favoring exploration early in evolution. Each mutation operator $\text{MUT}_{k}\in\{\text{MUT}_{k}\}_{k}$ corresponds to a KPI-aligned adjustment. Specifically, mutations include shifting tall items toward the center or lower layers to improve TallItem and CoG2D, filling small 2D voids below items to improve RelDen, compacting arrangements toward the origin to enhance AbsDen and RelDen, optimizing lateral contact to improve SideSup, and increasing bottom overlap to strengthen SurfSup. Additional diversification operations such as item, layer, or layer-addition swaps are also applied.